# Authorship Verification based on the Likelihood Ratio of Grammar Models

### v1

@@ Abstract
Authorship verification (AV) is the process of analyzing a set of documents to determine whether they were written by a specific author. This problem often arises in forensic scenarios, e.g., in cases where the documents in question constitute evidence for a crime. Existing state-of-the-art AV methods use computational solutions that are not supported by a plausible scientific explanation for their functioning and that are often difficult for analysts to interpret. To address this, we propose a method relying on calculating a quantity we call $\lambda_{G}$ (LambdaG): the ratio between the likelihood of a document given a model of the Grammar for the candidate author and the likelihood of the same document given a model of the Grammar for a reference population. These Grammar Models are estimated using $n$-gram language models that are trained solely on grammatical features. Despite not needing large amounts of data for training, LambdaG still outperforms other established AV methods with higher computational complexity, including a fine-tuned Siamese Transformer network. Our empirical evaluation based on four baseline methods applied to twelve datasets shows that LambdaG leads to better results in terms of both accuracy and AUC in eleven cases and in all twelve cases if considering only topic-agnostic methods. The algorithm is also highly robust to important variations in the genre of the reference population in many cross-genre comparisons. In addition to these properties, we demonstrate how LambdaG is easier to interpret than the current state-of-the-art. We argue that the advantage of LambdaG over other methods is due to fact that it is compatible with Cognitive Linguistic theories of language processing.
@@ 1 Introduction
Over the course of time, various research disciplines have emerged to deal with the analysis of the language of authors from different perspectives. This has led to an interdisciplinary branch of research that is often referred to as digital text forensics in the literature (e. g., [90, 10, 40, 95, 38]). Two well-known and closely related sub-areas of this branch of research are authorship attribution (AA) and authorship verification (AV). AA is concerned with the task of attributing an anonymous text to the most likely author based on a set of sample documents of known authors. AV, on the other hand, deals (in its simplest form) with the task of deciding whether two documents were written by the same author. Our focus in this paper is on AV, which represents a reformulation of AA, based on the fact that any AA problem can be decomposed into a series of AV problems [67].
@@ 1.1 Decision Problems
    $\textrm{AV}_{\,\textrm{Batch}}$: Given two sets of documents $\mathbb{D}_{\mathcal{A}}=\{\mathcal{D}^{\,(a)}_{1},\mathcal{D}^{\,(a)}_{2},\ldots\}$ and $\mathbb{D}_{\mathcal{B}}=\{\mathcal{D}^{\,(b)}_{1},\mathcal{D}^{\,(b)}_{2},\ldots\}$ each of which written by a single author: Are the documents in $\mathbb{D}_{\mathcal{A}}$ and $\mathbb{D}_{\mathcal{B}}$ written by the same author?
@@
    $\textrm{AV}_{\,\textrm{Known}}$: Given a set of documents $\mathbb{D}_{\mathcal{A}}=\{\mathcal{D}_{1},\mathcal{D}_{2},\ldots\}$ by the same author $\mathcal{A}$ as well as a document $\mathcal{D}_{\mathcal{U}}$ of an unknown author $\mathcal{U}$: Has $\mathcal{A}$ also written $\mathcal{D}_{\mathcal{U}}$ such that $\mathcal{A}=\mathcal{U}$ holds?
In a forensic context the process of assisting the trier-of-facts in finding answers to these problems is referred to as Forensic Text Comparison (FTC) [49]. The focus in this paper is on all three of these decisions problems. For the rest of this paper, we adopt the following notation: a document denoted by $\mathcal{D}_{\mathcal{U}}$, represents a document of an unknown author $\mathcal{U}$, whereas a document denoted by $\mathcal{D}_{\mathcal{A}}$ represents a document of a known author $\mathcal{A}$. In all decision problems, including $\textrm{AV}_{\,\textrm{Core}}$ where the denotation of which document is known or unknown can be arbitrary, an AV method seeks to answer the question of whether $\mathcal{A}=\mathcal{U}$ ($\mathcal{A}$ authored $\mathcal{D}_{\mathcal{U}}$) or $\mathcal{A}\neq\mathcal{U}$ ($\mathcal{A}$ did not author $\mathcal{D}_{\mathcal{U}}$) holds. In this context, a known verification case with a matching and non-matching authorship is denoted by Y- and N-case, respectively. Moreover, we denote $\neg\mathcal{A}$ as any author other than $\mathcal{A}$.
1.2 Applications of Authorship Verification
For all the mentioned forms of decision problems, numerous AV applications can be found across different domains. In forensic linguistics, for example, AV can be used to analyze malicious texts such as ransom notes, blackmail letters or other non-malicious texts that are used as evidence in investigatory, criminal or civil contexts. In the field of journalism and media, AV can play a vital role when it comes to confirming the authorship of news articles or opinion pieces or investigating cases of ghostwriting and unauthorized publication of content. In academia, AV can be used to analyze research articles, dissertations or other written content, for example, to resolve a plagiarism claim, or contract cheating from students, and thereby overall contribute to combating academic misconduct. In the context of social media, AV can also be employed to detect fake accounts or multiple accounts belonging to the same user. Moreover, as part of email and communication analysis, AV can be used to verify the authorship of emails in cases of fraud, harassment or cybercrime. Finally, in the Digital Humanities, AV can be used to resolve matters that deal with disputed historical evidence or to provide evidence in favour or against the origin of literary works.
@@ 1.3 Categories of AV Methods
According to Halvani et al. [43, 47], AV methods can be divided into three possible categories, namely unary, binary-intrinsic and binary-extrinsic AV methods [^1] A summarized visual representation of these three categories can be seen in [47, Figure 1]. All three rely on a decision criterion $\theta$, in order to accept or reject $\mathcal{A}$ as the author of $\mathcal{D}_{\mathcal{U}}$. The way in which $\theta$ is determined is what primarily distinguishes these three categories:
@@
  • Binary-intrinsic AV methods require a given training corpus, consisting of verification cases with a ground truth regarding the Y- (same-authorship) and N-classes (different-authorship), in order to determine $\theta$. Binary-intrinsic AV methods treat verification cases that include the unknown and known document(s) of $\mathcal{A}$, as a single unit $X$ (e. g., a feature vector). If $X$ is more similar to the Y-cases, the method accepts $\mathcal{A}$ as the author of $\mathcal{D}_{\mathcal{U}}$. On the other hand, if $X$ is more similar to the N-cases, $\mathcal{D}_{\mathcal{U}}$ is assumed to be written by another author. In any case, the decision is made solely on the basis of $X$ and the training classification model that incorporates $\theta$.
@@
  • Binary-extrinsic AV methods require external, so-called impostor [70] documents (which have not been written by $\mathcal{A}$) in order to determine $\theta$. To obtain impostor documents $\mathbb{I}=\{I_{1},I_{2},\ldots\}$, different strategies can be employed. Koppel and Winter [70], for example, used two approaches: (1) collecting documents using search engine queries and (2) using a fixed set of already crawled documents. Binary-extrinsic AV methods assume $\mathcal{D}_{\mathcal{U}}$ to be written by $\mathcal{A}$ only if $\mathcal{D}_{\mathcal{U}}$ is consistently more similar to $\mathcal{D}_{\mathcal{A}}$ (or alternatively, the documents in $\mathbb{D}_{\mathcal{A}}$) despite the presence of distractor impostor documents $\{I_{1},I_{2},\ldots\}$ after calculating a similarity score using a bootstrapping approach.
@@
In a forensic context, the adherence to the likelihood ratio paradigm for forensic science implies that an AV method does not return $\theta$ but instead returns a likelihood ratio that quantifies the strength of the evidence in favour or against the two propositions. In addition to the three types above, therefore, it is also possible to classify an AV method as forensic when the method returns a likelihood ratio.
@@
In a review of the field, Stamatatos [102] concluded that binary-extrinsic methods tend to outperform the other categories. In particular, a binary-extrinsic AV method that has been found to be very successful over several independent studies is the well-known Impostors Method (IM). The basic idea behind IM can be traced back to another successful AV method introduced by Koppel and Schler [65] and called Unmasking and later modified by Koppel et al. in [68]. The key insight of IM is to collect a set of texts that are plausible impostors for $\mathcal{D}_{\mathcal{A}}$ and then calculate a similarity score for $\mathcal{D}_{\mathcal{A}}$ and $\mathcal{D}_{\mathcal{U}}$ not once but a large number of times while withholding a random sub-set of the features at every iteration. The rationale is that only the real author of the disputed text is so similar to it to withstand random permutations of the features set, which, for IM, is the set of the most frequent character 4-grams. Koppel et al. [67] proposed that to choose the impostors the key principle should be to identify a relevant universe, choose linguistically similar texts, and then randomly sample from them.
@@
IM found immediate external validation in the PAN 2013 competition, where a variant called GenIM (General Impostors Method), proposed by Seidman [98], outperformed all other submitted AV methods. As suggested by Koppel et al. [66], GenIM made use of collected impostors document from the web. The method then used the relative frequency of single words weighted by their document frequency (a quantity usually called tf-idf) and the Min-Max (aka Ružička or fuzzy Jaccard) similarity. Successful results were then found by Koppel and Winter [70] for the $\textrm{AV}_{\,\textrm{Core}}$ decision problem, again using the Min-Max similarity between vectors of character 4-grams tf-idf. For the impostors, Koppel and Winter [70] tried various configurations, including with impostors generated on the fly using web searches or sampling from genre-compatible corpora. They discovered that genre compatibility permits the use of fewer impostors but that, equally, a large number of impostors ($\approx 200$) generated using web searches can compensates for lack of genre control. Their results demonstrated that it is possible to solve this challenging two-text only verification problem even for texts as short as 500 words.
@@
Although IM’s performance over the years is unquestionable, the four main drawbacks of the method are that it is not deterministic, that it suffers from a high run-time [41], that it relies on choosing impostors from the same genre, and that its results are often hard to interpret. The first issue is typically addressed by running several iterations and averaging the results but this additional procedure only leads to even longer run-times. The third issue is not strictly speaking an issue of IM but a well-known problem for both AV and AA, as virtually all methods to solve theses tasks tend to be highly affected by genre variation [5]. Regarding IM, Halvani [41] found that the performance significantly decreases when the set of impostors is chosen from another genre. For example, performing AV on a corpus of chat logs using a corpus of chat logs leads to a performance of 0.92 in accuracy. However, when using business emails or academic articles as impostors, this performance drops to almost chance level at 0.54. Only when the genres are relatively similar (e. g., chat logs compared to blogs) then, although a significant drop in performance is still evident, the results are still above chance (0.74).
@@
1.4 Effect of Content
This issue of the effect of content is one of a number of biases identified by Bevendorff et al. [11] on the basis of the 2014 and 2015 PAN competitions. These biases were taken in consideration by PAN starting from its 2020 edition [57]. However, the problem of using features that can potentially capture topic or content instead of authorship remains in many verification systems, including more recent ones.
One of the first approaches dealing with content bias in the context of authorship analysis was the topic-masking technique TextDistortion [103], which is simply based on replacing any word that is not among the top most frequent words in a reference corpus with an asterisk. TextDistortion was found to actually improve the results in cross-topic and cross-genre analyses. In 2021, Halvani and Graner [42] tested TextDistortion against another algorithm they introduced called POSNoise, which is a pre-processing method that consists in the replacement of words that are unlikely to be functional with their Part-Of-Speech (POS). They found that this method clearly improves the algorithms in cross-domain cases while often resulting in poorer performance in cases that are not cross-domain. These results are interpreted by them as evidence that this algorithm should be applied for forensic purposes because it is capable of removing the effect of content. Among the methods that benefited from this pre-processing was IM, which was also one of the most successful ones in their evaluations. A similar feature type to avoid content bias is the use of frame $n$-grams [82]. Similarly to POSNoise, this feature set consists of all function words and with all nouns, verbs, adjectives, adverbs, and pronouns being replaced with their corresponding POS.
@@ 1.4 Effect of Content
Another approach to counteract content bias, called TAVeer, was proposed by Halvani et al. [44]. In contrast to TextDistortion and POSNoise, TAVeer represents a stand-alone AV method based on ensembles of topic-agnostic (TA) feature categories (namely punctuation $n$-grams, TA sentence and sentence-initial tokens, TA sentence endings, TA token $n$-grams and TA masking token $n$-grams). The TA feature categories themselves cover a total of $\approx 1000$ words and phrases that originate from 20 linguistic categories such as punctuation marks, conjunctions, determiners, prepositions, contractions, auxiliary verbs, transition words and others, which are known in the literature [86, 12, 106] to be content and topic independent.
1.5 Language Models and Compression Methods

### v2

# Grammar as a Behavioral Biometric: Using Cognitively Motivated Grammar Models for Authorship Verification

@@ Abstract
Authorship Verification (AV) is a key area of research in digital text forensics, which addresses the fundamental question of whether two texts were written by the same person. Numerous computational approaches have been proposed over the last two decades in an attempt to address this challenge. However, existing AV methods often suffer from high complexity, low explainability and especially from a lack of clear scientific justification. We propose a simpler method based on modeling the grammar of an author following Cognitive Linguistics principles. These models are used to calculate $\lambda_{G}$ (LambdaG): the ratio of the likelihoods of a document given the candidate’s grammar versus given a reference population’s grammar. Our empirical evaluation, conducted on twelve datasets and compared against seven baseline methods, demonstrates that LambdaG achieves superior performance, including against several neural network-based AV methods. LambdaG is also robust to small variations in the composition of the reference population and provides interpretable visualizations, enhancing its explainability. We argue that its effectiveness is due to the method’s compatibility with Cognitive Linguistics theories predicting that a person’s grammar is a behavioral biometric.
@@ 1 Introduction
Over time, various research disciplines have emerged to deal with the analysis of the language of authors from different perspectives. This has led to an interdisciplinary branch of research that is often referred to as digital text forensics (e. g., [103, 14, 49, 109, 47]). Two well-known and closely related sub-areas of this branch of research are Authorship Attribution (AA) and Authorship Verification (AV). AA is concerned with the task of attributing an anonymous text to the most likely author based on a set of sample documents of known authors. AV, on the other hand, deals (in its simplest form) with the task of deciding whether two documents were written by the same author. In this paper, we focus on AV, which can be seen as a reformulation of AA, given that any AA problem can be decomposed into a series of AV problems [81].
@@ 1.1 Decision Problems
    $\textrm{AV}_{\,\textrm{Batch}}$: Given two sets of documents $\mathbb{D}_{\mathcal{A}}=\{\mathcal{D}^{\,\mathcal{A}}_{1},\mathcal{D}^{\,\mathcal{A}}_{2},\ldots\}$ and $\mathbb{D}_{\mathcal{B}}=\{\mathcal{D}^{\,\mathcal{B}}_{1},\mathcal{D}^{\,\mathcal{B}}_{2},\ldots\}$ each of which written by a single author: Were the documents in $\mathbb{D}_{\mathcal{A}}$ and $\mathbb{D}_{\mathcal{B}}$ written by the same author?
@@
    $\textrm{AV}_{\,\textrm{Known}}$: Given a set of documents $\mathbb{D}_{\mathcal{A}}=\{\mathcal{D}_{1},\mathcal{D}_{2},\ldots\}$ by the same author $\mathcal{A}$ as well as a document $\mathcal{D}_{\mathcal{U}}$ of an unknown author $\mathcal{U}$: Has $\mathcal{A}$ also written $\mathcal{D}_{\mathcal{U}}$ (i. e., $\mathcal{A}=\mathcal{U}$)?
In a forensic context the process of assisting the trier-of-facts in finding answers to these problems is referred to as Forensic Text Comparison (FTC) [62]. The focus in this paper is on all three of these decisions problems. For the rest of this paper, we adopt the following notation: a document denoted by $\mathcal{D}_{\mathcal{U}}$, represents a document of an unknown author $\mathcal{U}$, whereas a document denoted by $\mathcal{D}_{\mathcal{A}}$ represents a document of a known author $\mathcal{A}$. In all decision problems, including $\textrm{AV}_{\,\textrm{Core}}$ where the denotation of which document is known or unknown can be arbitrary, an AV method seeks to answer the question of whether $\mathcal{A}=\mathcal{U}$ ($\mathcal{A}$ authored $\mathcal{D}_{\mathcal{U}}$) or $\mathcal{A}\neq\mathcal{U}$ ($\mathcal{A}$ did not author $\mathcal{D}_{\mathcal{U}}$) holds. In this context, a known verification case with matching and non-matching authorship is denoted by Y- and N-case, respectively. Moreover, we denote $\neg\mathcal{A}$ as any author other than $\mathcal{A}$.
1.2 Applications of AV
AV can be used in many different areas and offers a wide range of possible applications. In forensic linguistics, AV is instrumental in analyzing texts related to criminal activity, such as ransom notes, blackmail letters, or other documents used as evidence in investigative, criminal, or civil cases. For example, Iqbal et al. [60] used a NIST-based speaker recognition system to verify the authorship of malicious emails in forensic contexts. Similarly, Remmide et al. [107] employed AV to detect phishing attacks. In journalism and media, AV plays a vital role in authenticating the authorship of news articles and opinion pieces, as well as detecting ghostwriting [4]. In academia, AV is crucial for detecting academic misconduct, such as contract cheating [122] and plagiarism [42]. Social media platforms also benefit from AV, especially when it comes to detecting fake accounts or multiple accounts operated by the same person. In this context, Igawa et al. [59] and Barbon et al. [67] demonstrated the effectiveness of AV regarding the identification of compromised accounts on X/Twitter. Lundmark et al. [87], on the other hand, used AV to link user accounts on discussion forums, whereas Weerasinghe et al. [134] applied AV to mitigate abuse in online communities. Finally, in the Digital Humanities, AV has also become an important instrument for clarifying questions about the authorship of historical documents and authenticating literary works. Kestemont et al. [72], for example, applied AV to authenticate historical writings attributed to Julius Caesar, while Tuccinardi [126] used AV methods to investigate the authenticity of Pliny the Younger’s letter to Emperor Trajan.
@@ 1.3 Categories of AV Methods
AV methods can be divided into three different categories: unary, binary-intrinsic and binary-extrinsic [54, 58]. All three rely on a decision criterion $\theta$, in order to accept or reject $\mathcal{A}$ as the author of $\mathcal{D}_{\mathcal{U}}$. The way in which $\theta$ is determined is what primarily distinguishes these three categories:
@@
  • Binary-intrinsic AV methods require a given training corpus, consisting of verification cases with a ground truth regarding the same-authorship (Y) and different-authorship (N) classes, in order to determine $\theta$. Binary-intrinsic AV methods treat verification cases that include the unknown and known document(s) of $\mathcal{A}$, as a single unit $X$ (e. g., a feature vector). If $X$ is more similar to the Y-cases, the method accepts $\mathcal{A}$ as the author of $\mathcal{D}_{\mathcal{U}}$. On the other hand, if $X$ is more similar to the N-cases, $\mathcal{D}_{\mathcal{U}}$ is assumed to be written by another author. In any case, the decision is made solely on the basis of $X$ and the training classification model that incorporates $\theta$.
@@
  • Binary-extrinsic AV methods require external, so-called impostor [84] documents (which have not been written by $\mathcal{A}$) in order to determine $\theta$. To obtain impostor documents $\mathbb{I}=\{I_{1},I_{2},\ldots\}$, different strategies can be employed. Koppel and Winter [84], for example, used two approaches: (1) collecting documents using search engine queries and (2) using a fixed set of already crawled documents. Binary-extrinsic AV methods assume $\mathcal{D}_{\mathcal{U}}$ to be written by $\mathcal{A}$ only if $\mathcal{D}_{\mathcal{U}}$ is consistently more similar to $\mathcal{D}_{\mathcal{A}}$ (or alternatively, the documents in $\mathbb{D}_{\mathcal{A}}$) than any of a number of distractor impostor documents $\{I_{1},I_{2},\ldots\}$ using a bootstrapping approach.
@@
In a forensic context, the adherence to the likelihood ratio paradigm for forensic science implies that an AV method does not return a Y- or N-decision but a likelihood ratio that quantifies the strength of the evidence in favor or against the two propositions. In addition to the three types above, therefore, it is also possible to classify an AV method as forensic when the method returns a likelihood ratio.
1.4 Existing AV Solutions
This section describes the advantages and limitations of some well-known existing AV methods that have demonstrated a consistent degree of effectiveness over time. This section covers in particular those methods that have found significant validation across the yearly competitions held at the PAN series of conferences focused on stylometry and digital text forensics. Since its launch in 2007, PAN has hosted 73 computational challenges, known as shared tasks, in areas such as authorship analysis, computational ethics, and evaluating the originality of writing, as well as curated 57 evaluation datasets, in addition to nine datasets contributed by the community [10].
1.4.1 The Impostors Method
@@
In a review of the field, Stamatatos [118] concluded that binary-extrinsic methods tend to outperform other types of methods. A widely recognized and successful binary extrinsic AV method that has been repeatedly validated in independent studies is the Impostors Method (IM), which has been proposed by Koppel and Winter [84]. The basic idea behind IM can be traced back to another successful AV method called Unmasking, initially introduced by Koppel and Schler [79] and later refined by Koppel et al. in [82]. Given a verification case $c=(\mathcal{D}_{\mathcal{U}},\mathcal{D}_{\mathcal{A}})$, the key insight of IM is to collect a set of texts that are plausible impostors for $\mathcal{D}_{\mathcal{A}}$ and then calculate a similarity score for $\mathcal{D}_{\mathcal{U}}$ and $\mathcal{D}_{\mathcal{A}}$ not once but a large number of times while withholding a random sub-set of the features at every iteration. The rationale is that only the real author of the disputed text $\mathcal{D}_{\mathcal{U}}$ is so similar to it to withstand random permutations of the features set, which, for IM, is the set of the most frequent character $4$-grams. Koppel et al. [81] suggested that the key principle to choose the impostors should be to identify a relevant population, select linguistically similar texts, and then randomly sample from it.
@@
IM found immediate external validation in the PAN 2013 competition, where a variant called GenIM (General Impostors Method), proposed by Seidman [113], outperformed all other submitted AV methods. As suggested by Koppel et al. [80], GenIM made use of collected impostors document from the web. The method then used the relative frequency of single words weighted by their document frequency (a quantity usually called tf-idf) and the Min-Max (aka Ružička or fuzzy Jaccard) similarity measure. Successful results were then found by Koppel and Winter [84] for the $\textrm{AV}_{\,\textrm{Core}}$ decision problem, again using Min-Max applied to vectors of character $4$-grams weighted by tf-idf. They explored various configurations for impostors, including dynamically generating them through web searches or sampling from genre-compatible corpora. Koppel and Winter discovered that genre compatibility permits the use of fewer impostors but that, equally, a large number of impostors ($\approx 200$) generated through web searches can compensate for the absence of genre control. Their results demonstrated that it is possible to solve this challenging two-text only verification problem even for texts as short as 500 words.
@@
Although IM’s performance over the years is unquestionable, the four main drawbacks of the method are that it is not deterministic, that it suffers from a high run-time [51], that it relies on choosing impostors from the same genre, and that its results are often hard to interpret. The first issue is typically addressed by running several iterations and averaging the results but this additional procedure only leads to even longer run-times. The third issue is not strictly speaking an issue of IM but a well-known problem for both AV and AA, as virtually all methods to solve theses tasks tend to be highly affected by genre variation [8]. For IM, Halvani [51] found that the performance (in terms of Accuracy [^1] Note that we use the capitalized form ”Accuracy” to clarify that we are talking about a specific machine learning evaluation metric, rather than the general concept of ”accuracy”.) declines significantly when impostors are drawn from a different genre. For example, performing AV on a corpus of chat logs achieves an Accuracy of 0.92 when using chat logs as impostors. However, when using business emails or academic articles as impostors, this performance drops to almost chance level at 0.54. Even with relatively similar genres, such as chat logs and blogs, the performance still decreases notably, though it remains above chance (0.74).
@@
1.4.2 Content-agnostic Approaches
This content effect limitation of IM is also one of a number of biases identified by Bevendorff et al. [14] on the basis of the 2014 and 2015 PAN competitions. These biases were taken in consideration by PAN starting from its 2020 edition [70]. However, the problem of using features that can potentially capture topic or content instead of authorship remains in many verification systems, including more recent ones.
One of the first approaches dealing with content bias in the context of authorship analysis was the topic-masking technique TextDistortion [119], which is simply based on replacing any word [^2] As an alternative, Stamatatos [119] proposed an additional variant of replacing not the entire word $w$ but instead each character in it with an asterisk. $w$ that is not among the top most frequent words in a reference corpus with an asterisk. TextDistortion was found to actually improve the results in cross-topic and cross-genre analyses. In 2021, Halvani and Graner [53] tested TextDistortion against another algorithm they introduced called POSNoise (see Section 6.2), which is a pre-processing method that consists in the replacement of words that are unlikely to be functional with their Part-Of-Speech (POS). They found that this method clearly improves the algorithms in cross-domain cases while often resulting in poorer performance in cases that are not cross-domain. These results are interpreted by them as evidence that this algorithm should be applied for forensic purposes because it is capable of removing the effect of content. Among the methods that benefited from this pre-processing was IM, which was also one of the most successful ones in their evaluations.
@@ 1.4.2 Content-agnostic Approaches
Another approach to counteract content bias, called TAVeer  [^3] TAVeer stands for Topic-agnostic Authorship Verifier based on equal error rate”., was proposed by Halvani et al. [55]. In contrast to TextDistortion and POSNoise, TAVeer represents a stand-alone AV method based on ensembles of topic-agnostic (TA) feature categories (namely punctuation $n$-grams, TA sentence and sentence-initial tokens, TA sentence endings, TA token $n$-grams and TA masked token $n$-grams). The TA feature categories themselves cover a total of $\approx 1000$ words and phrases that originate from 20 linguistic categories such as punctuation marks, conjunctions, determiners, prepositions, contractions, auxiliary verbs, transition words and others, which are known in the literature [100, 17, 124] to be content and topic independent. Given a verification case $c=(\mathcal{D}_{\mathcal{U}},\mathcal{D}_{\mathcal{A}})$ and a set of selected feature categories, TAVeer constructs two feature vectors for each category, which represent the $L^{1}$-normed occurrence frequencies of specific features (such as the conjunction word “and”) in the corresponding document. Afterwards, for each category, the Manhattan distance between two vectors is calculated and calibrated into the interval $[0,1]$, where $0.5$ represents a decision boundary. The median of the calibrated values over all selected categories, compared to $0.5$ then represents the final decision of $\mathcal{A}=\mathcal{U}$ or $\mathcal{A}\neq\mathcal{U}$. The thresholds and the optimal subset of selected TA feature categories are optimized, by selecting the ones that yield the best accuracy on a given training corpus.
1.4.3 Language Models and Compression Methods